Following information is available in respect of a company named, as Beta Ltd
Profit before taxesRs 50 lakhs, Dividend declared per share Rs2, Price of the share prevalent on stock exchange Rs 200, Applicable tax rate 35%,Share capital of company 5 lakh shares of Rs 10 each.
Calculate-1)Earnings per share 2) Dividend Pay-out Ratio 3) Price earnings ratio

Prepare the cash flow statement from investing activities of Alpha Creative Ltd for the year ended March31, 2019
1
Expert's answer
2019-11-14T08:52:58-0500

a. Earnings per share = Profit/Share capital,

"EPS = \frac{50×(1 - 0.35)lakhs}{5lakhs} = 6.5."

Dividend Pay-out Ratio = Dividend per share/Earnings per share,

"DP ratio = \frac{2}{6.5} = 0.31."

Price earnings ratio = Price per share/Earnings per share,

"PE ratio = \frac{200}{6.5} = 30.77."

b. Cash flow from investing activities is the section of a company's cash flow statement that displays how much money has been used in (or generated from) making investments during a specific time period.